Given f(x) = –x3 – 3x2 – 3x + 3, find f(–1), f(0), and f(3)
Answer by praseenakos@yahoo.com(507)

ANSWER:

Given f(x) = –x3 – 3x2 – 3x + 3, find f(–1), f(0), and f(3)

ANSWER:
Given f(x) = –x^3 – 3x^2 – 3x + 3,

To find f(–1), f(0), and f(3).....replace x by -1, 0 and 3 in the given equation.



f(x) = –x^3 – 3x^2 – 3x + 3,


f(-1) = –(-1)^3 – 3(-1)^2 – 3(-1) + 3


==> = -(-1) + 3 + 3 + 3



==> = 1 + 3 + 3 + 3

==> f(-1) = 10


f(0) = –(0)^3 – 3(0)^2 – 3(0) + 3


==> = -0 -0 -0 + 3



f(0) = 3



f(3) = –3^3 – 3* 3^2 – 3*3 + 3


==> = -27 - 3 * 9 - 9 + 3


==> = -27 -27 - 9 + 3


==> = -63 + 3


==> = -60





Hope you found the method useful.


Regards.


Praseena.